OpenStudy (anonymous):

f(x)=(1+16x^2)tan^-1(4x)...find the derivative...i got 128x/(1+16x^2)...its wrong...i keep missing a step

OpenStudy (vijay):

\[32x \tan^{-1} (4x) + 4\]

OpenStudy (anonymous):

how did u get that

OpenStudy (vijay):

differentiation by parts

OpenStudy (anonymous):

but the derivative of tan^-1 is 1/1+x^2 and you plug 4x into for the x....anc then continue...what happened to that equation

OpenStudy (vijay):

tan^-1(x) = 1/ 1 +x^2, but tan^-1(4x) = 4/(4^2 + x^2)

OpenStudy (anonymous):

ok so how did u get from that to the answer u gave

OpenStudy (anonymous):

i dont understand why you would use the derivative equation for one problem and not the other